Lars:

Craig solves the following system of equations using the elimination method. 2x+4y=16 4x+8y=−16 He chooses to eliminate the variable y. Which statement correct describes his process and solution? A) Craig multiplies 2x + 4y = 16 by 2 and then adds the equations. He finds the result is a contradiction, so there is no solution. B) Craig multiplies 2x + 4y = 16 by 2 and then adds the equations. He finds the result is an identity, so there are infinitely many solutions. C) Craig multiplies 2x + 4y = 16 by 2 and then subtracts the equations. He finds the result is an identity, so there are infinitely many solutions. D) Craig multiplies 2x + 4y = 16 by 2 and then subtracts the equations. He finds the result is a contradiction, so there is no solution.

yourdootcome:

ok so try multiplying the first equation by 2 what do you get?

yourdootcome:

you get 4x+8y=16 and see how the other one is 4x+8y=-16 that means there is no solution because both contradict each other

Lars:

Ohh okay. So Craig is subtracting then

Lars:

Or I think-

yourdootcome:

no he multiplies the first equation by 2 2(2x+4y)=(16)*2 = 4x+8y=32

yourdootcome:

and it says that 4x+8y=-16 on the second equation which directly contradicts the first equation which means the answer is D because 32 does not equal -16

Lars:

Ohh okay that clears up a bit of my confusion. Thank you